OpenStudy (anonymous):

Consider the functions f(x) = 3x2 and g(x) = 2(x + 9). Find f[g(−7)]. Please show the work? Thanks.

OpenStudy (anonymous):

Find g(-7) by plugging -7 in for g(x)

OpenStudy (anonymous):

\[g(-7)=2((-7)+9)\]

OpenStudy (anonymous):

\[g(-7)=4\]

OpenStudy (anonymous):

Then, plug 4 into f(x).

OpenStudy (anonymous):

\[f(4)=3(4)^{2}\]

OpenStudy (anonymous):

\[f(4)=48\]

OpenStudy (anonymous):

Plug in g(-7) into g(x) and then plug in that solution to f(x):\[g(-7)=2(-7+9)\]\[g(-7)=2(2)\]\[g(-7)=4\]now plug in \(4\) to \[f(x)=3(4)^2\]\[f(x)=3(16)\]\[f(x)=48\] and thus: \(\largef[g(-7)]=48\)

OpenStudy (anonymous):

Oh, thank you! I hate complicated functions like these. Thanks to both of you!

OpenStudy (anonymous):

No Problem :)

OpenStudy (anonymous):

Np..